Three resistors of 6-ohm resistance are connected in delta. Inside the delta, another 6-ohm resistor  is connected in wye. Find the resistance between any two corners. 




1
Expert's answer
2021-10-04T19:38:49-0400

"R_1=R+R=2R\Omega\\R_2=R+R=2R\Omega"

"R_1,R_2,R'" Is parallel combination

"\frac{1}{R_{net}}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R'}"

"\frac{1}{R_{net}}=\frac{1}{2R}+\frac{1}{2R}+\frac{1}{R}"

"\frac{1}{R_{net}}=\frac{4}{2R}\\R_{net}=\frac{R}{2}"